# Self-consistent study of topological superconductivity in two-dimensional quasicrystals

### v1

@@ Abstract
We study two-dimensional $s$-wave topological superconductivity with Rashba spin-orbit coupling and Zeeman field in Penrose and Ammann-Beenker quasicrystals. By solving the Bogoliubov-de Gennes equations self-consistently for not only the superconducting order parameter, but also the spin-dependent Hartree potential, we show the stable occurrence of TSC with broken time-reversal symmetry in both Penrose and Ammann-Beenker quasicrystals. The topological nature of the quasicrystalline system is signified by the Bott index $B$. Topological phase transitions are found to occur, where $B$ changes between 0 and ±$\pm 1$, as the chemical potential or Zeeman field is varied. In terms of self-consistent solutions, we demonstrate the existence of a Majorana zero mode per edge or vortex when $B=\pm 1$, consistently with the bulk-edge/defect correspondence for periodic systems.
@@ I INTRODUCTION
There are growing appeals for quasicrystals (QCs), unconventional materials with quasiperiodic order and rotational symmetry that is forbidden by crystallography [1], from the fundamental as well as application points of view. Mathematically, a QC can be obtained as projection of a periodic lattice in higher-dimensional space (‘hyperspace’ or superspace) called a hypercubic lattice onto lower dimensions. It is this relation of quasiperiodicity to higher dimensions that makes QCs all the more interesting, with additional degrees of freedom such as phasons and capability of hosting topological phenomena [2, 3, 4, 5]. While research on QCs has a long history since its first discovery in 1984 [6, 7, 8], QCs have recently been studied intensively [9, 10, 11, 12, 13]. QCs have been realised not only in solid-state materials, but also in optical systems [14, 15, 16] and quasicrystalline monolayers assembled on substrates [17, 18, 19, 20, 21, 22]. Superconductivity has been discovered in an icosahedral Al-Zn-Mg quasicrystal [23] and a van der Waals-layered Ta-Te quasicrystal [24]. Theoretically, conventional [25] as well as unconventional [26, 27] superconductivity in QCs has been discussed. Moreover, recent theoretical studies have focused on QCs as topological materials. For example, the integer quantum Hall effect [28, 29, 30], the quantum spin Hall effect [31, 32, 33], and higher-order topological phases [34, 35] have been explored. It has also been shown that QCs can host weak topological superconductivity phases, which can be characterised by a real-space weak topological invariant [36].
@@
The paper is organised as follows. In Sec. II Penrose and AB QCs and their representation in the perpendicular space – the complementary space to 2D real space – are introduced, and the BdG equations on the tight-binding model [48, 49, 58] and calculation of the Bott index are described briefly. Results of our self-consistent calculations are presented and discussed for Penrose and AB QCs in Sec. III.1 and III.2, respectively. The work is summarised and discussed further in Sec. IV.
@@ III.1.1 Topological phase transitions
In Ref. [58], by assuming a constant superconducting order parameter $\Delta_{i}\equiv\Delta$, topological phase transitions were found to occur as a function of the chemical potential $\mu$ near the lower band edge for fixed $\alpha$ and $h$ in both Penrose and AB QCs. These phase transitions follow the same condition for closing of the bulk spectral gap, $(W+\mu)^{2}+\Delta^{2}=h^{2}$, as for a square lattice [49], where half of the kinetic-energy bandwidth $W$ is obtained numerically. When it is solved for self-consistently, the order parameter $\Delta_{i}$ is not uniform and thus, there is no such gap-closing condition. Moreover, $\Delta_{i}$ acquires a non-negligible imaginary part at each lattice site as the inhomogeneity of the quasicrystalline lattice has similar effects as nonmagnetic impurities on the underlying $p\pm\imath p$ nature of TSC [71]. Here we show in terms of self-consistent solutions the occurrence of a topological phase transition as $\tilde{\mu}$ is varied. Figure 2 presents the lowest absolute value of the quasiparticle excitation energy $|\epsilon|_{{\rm{min}}}$ as a function of $\tilde{\mu}/t$ for a 3571-site Penrose QC for $h=-t$, $\alpha=t$, and $U=-5.2t$ with PBC. The spectral gap closes at $\tilde{\mu}\simeq 3.7625t$, where the Bott index changes from $B=0$ to $B=1$. For all values of $\tilde{\mu}$ studied here, the converged order parameter $\Delta_{i}$ is nonzero at all sites and the real part of the site average, $\bar{\Delta}=(1/N)\sum_{i}\Delta_{i}$, decreases monotonically from ${\rm Re}\bar{\Delta}\simeq 0.8t$ at $\tilde{\mu}=3.2t$ to ${\rm Re}\bar{\Delta}\simeq 0.3t$ at $\tilde{\mu}=4t$.
@@
By varying the Zeeman energy $h$ for fixed $\alpha$ and $\tilde{\mu}$, we also find a topological phase transition in our self-consistent solutions. This is illustrated for a 3571-site Penrose QC for $\alpha=t$, $\tilde{\mu}=-3.5t$, and $U=-5.5t$ with PBC in Fig. 3, where the lowest absolute excitation energy $|\epsilon|_{{\rm{min}}}$ is plotted as a function of $h/t$. The spectral gap closes at $h\simeq 1.3845t$, where the Bott index changes from 0 to $-1$. As $h$ varies from $h=t$ to $h=1.8t$, the effective Zeeman energy $\tilde{h}$ increases from $\tilde{h}\simeq 0.86t$ to $\tilde{h}\simeq 1.35t$ and the real part of the site-averaged order parameter monotonically decreases from ${\rm Re}\bar{\Delta}\simeq 0.8t$ to ${\rm Re}\bar{\Delta}\simeq 0.1t$, while $\Delta_{i}$ remains nonzero at all sites. It is interesting to note that the overall behaviour of the spectral gap as a function or $h$ or $\tilde{h}$ (not shown) around the topological phase transition is similar to the analytical formula for a square lattice [49] with a constant order parameter [72].
@@ IV CONCLUSION
In this work, we have focused on the kinetic-energy range near the top or bottom of the band, where the dispersion in Penrose and AB QCs is similar to that in a square lattice [58]. In contrast, the band structure can be very different from that of a square lattice in other energy ranges, e.g., due to possible formation of a pseudogap [74, 22, 25] and the presence of strictly localised states [75, 76, 77] in both QCs and with insulating gaps in Penrose QCs [78, 79]. It is thus interesting to explore different regions of the kinetic-energy band for the possible existence of TSC. Momentum-space analysis of the band structure in terms of the spectral function [25] will be useful for understanding the nature of TSC and topological phase transitions [80]. Such studies are currently under way.
@@ Acknowledgements.
We thank Rasoul Ghadimi, Aksel Kobiałka, Yoshi Maeno, Yuki Nagai, Ryo Okugawa, and Andrzej Ptok for stimulating discussions. K.T. is grateful to the Pauli Center for Theoretical Studies for support and the University of Zürich for hospitality, where part of the work was performed. This work was supported by the Natural Sciences and Engineering Research Council of Canada, JST SPRING (Grant No. JPMJSP2151), and the Japan Society for the Promotion of Science, KAKENHI (Grant No. JP19H05821). We also acknowledge support by the Tokyo University of Science Grant for International Joint Research. This research was made feasible in part by support provided by Calcul Québec, Compute Ontario, and the Digital Research Alliance of Canada (alliancecan.ca).

### v2

@@ Abstract
We study two-dimensional $s$-wave topological superconductivity with Rashba spin-orbit coupling and Zeeman field in Penrose and Ammann-Beenker quasicrystals. By solving the Bogoliubov-de Gennes equations self-consistently for not only the superconducting order parameter, but also the spin-dependent Hartree potential, we show the stable occurrence of topological superconductivity with broken time-reversal symmetry in both Penrose and Ammann-Beenker quasicrystals. The topological nature of the quasicrystalline system is signified by the Bott index $B$. Topological phase transitions are found to occur, where $B$ changes between 0 and ±$\pm 1$, as the chemical potential or Zeeman field is varied. In terms of self-consistent solutions, we demonstrate the existence of a Majorana zero mode per edge or vortex when $B=\pm 1$, consistently with the bulk-edge/defect correspondence for periodic systems.
@@ I INTRODUCTION
There are growing appeals for quasicrystals (QCs), unconventional materials with quasiperiodic order and rotational symmetry that is forbidden by crystallography [1], from the fundamental as well as application points of view. Mathematically, a QC can be obtained as projection of a periodic lattice in higher-dimensional space (‘hyperspace’ or superspace) called a hypercubic lattice onto lower dimensions. It is this relation of quasiperiodicity to higher dimensions that makes QCs all the more interesting, with additional degrees of freedom such as phasons and capability of hosting topological phenomena [2, 3, 4, 5]. While research on QCs has a long history since its first discovery in 1984 [6, 7, 8], QCs have recently been studied intensively [9, 10, 11, 12, 13]. QCs have been realised not only in solid-state materials, but also in optical systems [14, 15, 16] and quasicrystalline monolayers assembled on substrates [17, 18, 19, 20, 21, 22]. Superconductivity has been discovered in an icosahedral Al-Zn-Mg quasicrystal [23] and a van der Waals-layered Ta-Te quasicrystal [24]. Theoretically, unconventional features of superconductivity in QCs have been discussed in terms of the attractive Hubbard model [25, 26, 27] and its extension [28]. Moreover, recent theoretical studies have focused on QCs as topological materials. For example, the integer quantum Hall effect [29, 30, 31], the quantum spin Hall effect [32, 33, 34], and higher-order topological phases [35, 36] have been explored. It has also been shown that QCs can host weak topological superconductivity phases, which can be characterised by a real-space weak topological invariant [37].
@@
The paper is organised as follows. In Sec. II Penrose and AB QCs and their representation in the perpendicular space – the complementary space to 2D real space – are introduced, and the BdG equations on the tight-binding model [49, 50, 59] and calculation of the Bott index are described briefly. Results of our self-consistent calculations are presented and discussed for Penrose and AB QCs in Sec. III.1 and III.2, respectively. In Sec. III.3 the effects of randomly distributed nonmagnetic impurities on TSC in QCs are studied. The work is summarised and discussed further in Sec. IV. In Appendix, topological phase transitions in a square lattice are illustrated in terms of self-consistent solutions for comparison with the analogous results for QCs.
@@ III.1.1 Topological phase transitions
In Ref. [59], by assuming a constant superconducting order parameter $\Delta_{i}\equiv\Delta$, topological phase transitions were found to occur as a function of the chemical potential $\mu$ near the lower band edge for fixed $\alpha$ and $h$ in both Penrose and AB QCs. These phase transitions follow the same condition for closing of the bulk spectral gap, $(W+\mu)^{2}+\Delta^{2}=h^{2}$, as for a square lattice [50], where half of the kinetic-energy bandwidth $W$ is obtained numerically. When it is solved for self-consistently, the order parameter $\Delta_{i}$ is not uniform and thus, there is no such gap-closing condition. Moreover, $\Delta_{i}$ acquires a non-negligible imaginary part at each lattice site as the inhomogeneity of the quasicrystalline lattice has similar effects as nonmagnetic impurities on the underlying $p\pm\imath p$ nature of TSC [72]. Here we show in terms of self-consistent solutions the occurrence of a topological phase transition as $\tilde{\mu}$ is varied. Figure 2 presents the lowest absolute value of the quasiparticle excitation energy $|\epsilon|_{{\rm{min}}}$ as a function of $\tilde{\mu}/t$ for a 3571-site Penrose QC for $h=-t$, $\alpha=t$, and $U=-5.2t$ with PBC. The spectral gap closes at $\tilde{\mu}\simeq 3.7625t$, where the Bott index changes from $B=0$ to $B=1$. For all values of $\tilde{\mu}$ studied here, the converged order parameter $\Delta_{i}$ is nonzero at all sites and the real part of the site average, $\bar{\Delta}=(1/N)\sum_{i}\Delta_{i}$, decreases monotonically from ${\rm Re}\bar{\Delta}\simeq 0.8t$ at $\tilde{\mu}=3.2t$ to ${\rm Re}\bar{\Delta}\simeq 0.3t$ at $\tilde{\mu}=4t$. Comparing Fig. 2 with Fig. 13 for a square lattice with comparable $N$, it can be seen that $\tilde{\mu}$ at which the topological phase transition occurs is pushed up to a higher value in the Penrose QC. This makes sense from the fact that the kinetic-energy levels are more spread out in Penrose (as well as AB) QCs and go outside of the range $[-4t,4t]$ as for a square lattice [59].
@@
By varying the Zeeman energy $h$ for fixed $\alpha$ and $\tilde{\mu}$, we also find a topological phase transition in our self-consistent solutions. This is illustrated for a 3571-site Penrose QC for $\alpha=t$, $\tilde{\mu}=-3.5t$, and $U=-5.5t$ with PBC in Fig. 3, where the lowest absolute excitation energy $|\epsilon|_{{\rm{min}}}$ is plotted as a function of $h/t$. The spectral gap closes at $h\simeq 1.3845t$, where the Bott index changes from 0 to $-1$. As $h$ varies from $h=t$ to $h=1.8t$, the effective Zeeman energy $\tilde{h}$ increases from $\tilde{h}\simeq 0.86t$ to $\tilde{h}\simeq 1.35t$ and the real part of the site-averaged order parameter monotonically decreases from ${\rm Re}\bar{\Delta}\simeq 0.8t$ to ${\rm Re}\bar{\Delta}\simeq 0.1t$, while $\Delta_{i}$ remains nonzero at all sites. The overall behaviour of the spectral gap as a function or $h$ or $\tilde{h}$ (not shown) around the phase transition and in the topological region with $B=-1$ is similar to that in a square lattice, as can be seen in Fig. 14. However, compared to the square lattice, stronger Zeeman field is required for the transition to occur, and the spectral gap in the $B=-1$ region is significantly smaller in the QC and closes at $h\simeq 1.87t$. The TSC phase ceases to exist with vanishing order parameter ($|\Delta_{i}|\lesssim 10^{-5}t$) at all sites for $h\simeq 1.88t$ in the 3571-site Penrose QC.
@@ III.2.2 Majorana vortex bound state
III.3 Effects of nonmagnetic impurities
Due to its $p\pm\imath p$ nature, TSC tends to be affected readily by nonmagnetic impurities [74], which can be incorporated as the single-particle potential $v_{i}=V_{\rm imp}$ at a given lattice site $i$, namely, a local shift of the chemical potential by $-V_{\rm imp}$. For a square lattice, it has been shown that a single nonmagnetic impurity can cause spin-polarised midgap excitation and even phase transitions of the ground state as the impurity potential $V_{\rm imp}$ is varied [72]. Furthermore, in terms of the self-consistently solved, average impurity self-energy, the work of Ref. [75] shows that as the concentration of nonmagnetic impurities increases, the spectral gap and the superconducting transition temperature are reduced. And the stronger the Zeeman field, the more drastic these effects of nonmagnetic impurities are on TSC [72, 75].
What happens to TSC in QCs if nonmagnetic impurities are added randomly to the quasicrystalline lattice? We have examined the effects of nonmagnetic impurities with a constant potential $V_{\rm imp}$ distributed at $n_{\rm imp}$ random sites in both a square lattice and QCs. In Fig. 11 we present the lowest absolute value of the quasiparticle excitation energy $|\epsilon|_{{\rm{min}}}$ as a function of $n_{\rm imp}$ with $V_{\rm imp}=-t$ (orange circles) and $V_{\rm imp}=t$ (red and blue squares) in a $3571$-site Penrose QC with PBC for $h=-t$, $\alpha=t$, $\tilde{\mu}=3.8t$, and $U=-5.2t$. For both $V_{\rm imp}=\pm t$, the same set of random sites $\{i\}$ was used to place the nonzero $v_{i}$ for a given $n_{\rm imp}$. For the repulsive potential $V_{\rm imp}=t$, the spectral gap monotonically decreases as $n_{\rm imp}$ increases, showing that TSC is suppressed by nonmagnetic impurities as expected. At $n_{\rm imp}=103$ the gap closes and the Bott index changes from 1 to 0; and as $n_{\rm imp}$ is increased further, the spectral gap increases in the trivial superconducting phase with $B=0$.
Figure 11: Lowest absolute energy eigenvalue $|\epsilon|_{{\rm{min}}}/t$ as a function of the number $n_{\rm imp}$ of randomly distributed nonmagnetic impurities with $V_{\rm imp}=-t$ (orange circles) or $V_{\rm imp}=t$ (red and blue squares) in a Penrose QC with $N=3571$, $h=-t$, $\alpha=t$, $\tilde{\mu}=3.8t$, $U=-5.2t$, and PBC.
For the attractive potential $V_{\rm imp}=-t$, we have found the opposite effect that the spectral gap is enhanced as the impurities are added randomly to the QC, as can be seen in Fig. 11. The Bott index remains to be 1 and thus, the TSC phase is made more robust by randomly distributed nonmagnetic impurities. This may be reminiscent of a topological Anderson insulator [76] as a result of a topological phase transition from a normal metal caused by disorder (typically a random shift of the chemical potential at each lattice site), with increasing energy gap as the disorder is made stronger. In the current case, however, the increasing energy gap is basically due to the effective chemical potential of the system being increased as $n_{\rm imp}$ increases (see Fig. 2). In fact, as $n_{\rm imp}$ is increased further, the spectral gap reaches its maximum value $\simeq 0.18t$ at $n_{\rm imp}\approx 1200$, beyond which it decreases gradually and is about $0.08t$ at $n_{\rm imp}=3,000$. This is also consistent with the effective chemical potential of the system increasing, as such dependence of the gap on $\tilde{\mu}$ is expected for $\tilde{\mu}>4t$ by comparing Fig. 2 with Fig. 13.
A larger spectral gap implies a shorter effective coherence length. This is demonstrated in Fig. 12 for the same system presented in Fig. 8, a 9349-site Penrose QC with OBC and PBC in the $x$ and $y$ directions, respectively, for $h=t$, $\alpha=t$, $U=-5.2t$, and $\tilde{\mu}=3.8t$, but with $n_{\rm imp}=4,000$ and $V_{\rm imp}=-t$. The energy of the Majorana edge modes is $\epsilon\simeq 5\times 10^{-4}t$ and $3\times 10^{-10}t$, respectively, without and with the nonmagnetic impurities. Comparing Figs. 8 and 12, it is clear that the Majorana wave functions are confined much more strongly along and near each edge, and therefore, the two Majorana bound states at the two edges have negligible overlap.
Figure 12: (a,b) Electron and (c,d) hole amplitudes for spin (a,c) up and (b,d) down for the quasiparticle excitation with energy $\epsilon\simeq 10^{-10}t$ in the same system as in Fig. 8, but with $4,000$ randomly distributed nonmagnetic impurities with $V_{\rm imp}=-t$.
We have also performed calculation for $h=1.5t$, $\alpha=t$, $\tilde{\mu}=-3.5t$, and $U=-5.5t$ in a $3571$-site Penrose QC with PBC. Compared with the system discussed above, the spectral gap in this system with no impurities is $0.127t$, which is much larger than $0.05t$ at $n_{\rm imp}=0$ in Fig. 11 (see also Figs. 2 and 3), and the signs of $h$ and $\tilde{\mu}$ are opposite. When the sign of the chemical potential is changed, the effects of the opposite signs of $V_{\rm imp}$ are interchanged [72], so that the attractive potential only has the detrimental effect. For $V_{\rm imp}=-t$, because of the larger gap in the current system, the TSC phase persists up to $n_{\rm imp}=431$ before turning into the trivial superconducting phase. At the same time, as $|h|/t$ is larger, the effects of the nonmagnetic impurities are more drastic in that at $n_{\rm imp}=100$ the spectral gap is already reduced to $0.016t$. Likewise, for $V_{\rm imp}=t$, the spectral gap increases only slightly up to the maximum value $\simeq 0.135t$ at $n_{\rm imp}=145$ and then decreases though gradually to $0.055t$ at $n_{\rm imp}=3,000$. When the Zeeman energy is increased to $h=1.7t$, the spectral gap ($0.04t$ at $n_{\rm imp}=0$) starts to decrease monotonically with increasing $n_{\rm imp}\gtrsim 10$. Therefore, the enhancement of TSC by randomly distributed nonmagnetic impurities can occur only when the Zeeman field is not too strong.
Finally, we have performed analogous calculation on a $60\times 60$ square lattice with PBC for the two parameter sets used above, with $V_{\rm imp}=t$ or $-t$. In contrast to the results discussed above, for both parameter sets and for either sign of $V_{\rm imp}$, we have only found the detrimental effect on TSC by nonmagnetic impurities placed at random sites in a square lattice. The spectral gap starts to decrease as soon as $n_{\rm imp}$ is nonzero and continues to decrease as $n_{\rm imp}$ increases, though more slowly for $V_{\rm imp}=-t$ and $V_{\rm imp}=t$, respectively, for the former and latter parameter sets. For the former parameter set, the reduction of the gap can be understood again as due to a shift in the effective chemical potential of the system by looking at Fig. 13: Without any impurity, the spectral gap is maximum around $\tilde{\mu}=3.8t$ and shifting $\tilde{\mu}$ in either direction would decrease the gap. For the latter parameter set, even though the spectral gap in the clean system is twice as large as that in the 3571-site Penrose QC for $h=1.5t$ (see Fig. 14), the relatively large Zeeman energy presumably dominates the effects of the nonmagnetic impurities.
@@ IV CONCLUSION
Throughout this work, we have used the vertex model where the electron hops from site to site along the edge of a tile constituting the quasiperiodic tiling and as all edges are of the same length, this makes both Penrose and AB QCs bipartite. This may not be realistic in modelling real systems in that the tile edge is not necessarily the shortest link that connects two lattice sites. However, in the work of Ref. [59] it has been shown for a uniform order parameter that adding shortest-distance links, thus breaking the bipartiteness, does not change the results in any fundamental way. We presume that this is the case also when the mean fields are solved self-consistently.
We have also neglected the orbital effect of the magnetic field perpendicular to the quasicrystal. In this work, we have focused on the kinetic-energy range near the top or bottom of the band, where the dispersion in Penrose and AB QCs can be regarded quadratic as in a square lattice [59]. Assuming a quadratic dispersion, the effective mass and hence a uniform magnetic field that corresponds to the Zeeman energy $h/t$ can be estimated. By incorporating the magnetic field in terms of the Peierls factor [58] and using OBC, we have checked for some of the systems studied above that even though it tends to require more iterations, the converged mean fields are very similar to those without the Peierls factor. Moreover, we have confirmed the appearance of Majorana zero modes at the vortex centre and along the edges in the system presented in Fig. 10 with the Peierls factor included, with the excitation energy $\epsilon\simeq 4\times 10^{-5}t$.
In contrast to the kinetic-energy range that we have focused on in this work, the band structure can be very different from that of a square lattice in other energy ranges, e.g., due to possible formation of a pseudogap [77, 22, 26] and the presence of strictly localised states [78, 79, 80] in both QCs and with insulating gaps in Penrose QCs [81, 82]. It is thus interesting to explore different regions of the kinetic-energy band for the possible existence of TSC [83]. Momentum-space analysis of the band structure in terms of the spectral function [26, 59, 84] will be useful for understanding the nature of TSC and topological phase transitions. Such studies are currently under way.
As an experimental setup for creating a two-dimensional quasicrystalline topological superconductor, we propose a Pb monolayer with inherent $s$-wave superconductivity and strong Rashba spin-orbit coupling deposited [19] on the fivefold-symmetric surface of a three-dimensional quasicrystal such as Al-Pd-Mn [17, 19, 20] or Al-Cu-Fe [18], with a single-layer of Co [48] deposited on top for generating the Zeeman field. Moreover, superconductivity has recently been observed in a van der Waals-layered Ta1.6Te quasicrystal [24], which consists of layered two-dimensional dodecagonal quasicrystals [85], with evidence of strong spin-orbit coupling [86]. Therefore, Ta1.6Te quasicrystals doped with magnetic impurities or layered with a ferromagnetic insulator are potential candidate materials for realising TSC with broken time-reversal symmetry.
@@ Acknowledgements.
We thank Rasoul Ghadimi, Aksel Kobiałka, Yoshi Maeno, Yuki Nagai, Ryo Okugawa, and Andrzej Ptok for stimulating discussions. K.T. is grateful to the Pauli Center for Theoretical Studies for support and the University of Zürich for support and hospitality, where part of the work was performed. This work was supported by the Natural Sciences and Engineering Research Council of Canada, JST SPRING (Grant No. JPMJSP2151), and the Japan Society for the Promotion of Science, KAKENHI (Grant No. JP19H05821). We also acknowledge support by the Tokyo University of Science Grant for International Joint Research. This research was made feasible in part by support provided by Calcul Québec, Compute Ontario, and the Digital Research Alliance of Canada (alliancecan.ca).
Appendix A Topological phase transitions in periodic systems
Figure 13: Lowest absolute energy eigenvalue $|\epsilon|_{{\rm{min}}}/t$ as a function of the chemical potential $\tilde{\mu}/t$ in a $60\times 60$ square lattice for $h=-t$, $\alpha=t$, $U=-5.2t$, and PBC.
Figure 14: Lowest absolute energy eigenvalue $|\epsilon|_{{\rm{min}}}/t$ as a function of the Zeeman energy $h/t$ in a $60\times 60$ square lattice for $\alpha=t$, $\tilde{\mu}=-3.5t$, $U=-5.5t$, and PBC.
Here we illustrate topological phase transitions in periodic systems which are analogous to those presented in Figs. 2 and 3 for $3571$-site Penrose QCs. In Fig. 13 the lowest absolute excitation energy $|\epsilon|_{{\rm{min}}}$ is plotted as a function of $\tilde{\mu}/t$ for a $60\times 60$ square lattice with PBC for $h=-t$, $\alpha=t$, and $U=-5.2t$. The converged order parameter $\Delta_{0}$, which is uniform and real, decreases monotonically from $0.86t$ at $\tilde{\mu}=3.2t$ to $0.32t$ at $\tilde{\mu}=4t$ (not shown). At $\tilde{\mu}\simeq 3.51t$, $\Delta_{0}\simeq 0.61t$ and the effective Zeeman energy $\tilde{h}\simeq-0.78t$, satisfying the gap closing condition for $\tilde{\mu}>2t$ [50], $(4t-\tilde{\mu})^{2}+\Delta_{0}^{2}=\tilde{h}^{2}$, and the TKNN number [52] changes from 0 to $-1$. The Bott index ($B$) as defined in Eq. (9) has the opposite sign and thus changes from 0 to 1.
Shown in Fig. 14 is the absolute minimum eigenvalue $|\epsilon|_{{\rm{min}}}$ as a function of the Zeeman energy $h/t$ for a $60\times 60$ square lattice with PBC for $\alpha=t$, $\tilde{\mu}/t=-3.5t$, and $U=-5.5t$. From $h=t$ to $h=1.8t$, $\Delta_{0}$ ($\tilde{h}$) decreases (increases) monotonically from $0.85t$ ($0.82t$) to $0.22t$ ($1.33t$). At $h\simeq 1.115t$, $\tilde{h}\simeq 0.88t$ and $\Delta_{0}\simeq 0.72t$, where the condition for $\tilde{\mu}\leq-2t$ for a topological phase transition [50], $(4t+\tilde{\mu})^{2}+\Delta_{0}^{2}=\tilde{h}^{2}$, is satisfied and the TKNN number ($B$) changes from 0 to 1 ($-1$).